# Applications to Monolayers, Nanotubes and Nanotori

### v1

@@ III.2 Elastic Energy of single layer hBN Nanotubes and the Bending Rigidity of hBN Monolayers
which is equivalent to the common expression $\sqrt{3}l_{\mathrm{BN}}\left(n^{2}+nm+m^{2}\right)$ [72] involving pairs $(n,m)$ of chiral indices, upon setting $(n,m)=(n_{\mathrm{ZZ}},0)$ for ZZ and $n=m=n_{\mathrm{AC}}$ for AC nanotubes. The initial length of the hBN nanotube is
@@
    The inner level entails minimizing the potential energy of a NT with a fixed (input) length and returning the stress along the axial direction (output). The minimization is performed sequentially using the Polak-Ribiere conjugate-gradient [73] and the Hessian-free truncated Newton algorithms [64] until the norm of the global force vector is less than $10^{-9}\textrm{ ev}/$Å.
@@
    In the outer level the length of the NT is optimized until the stress of the minimized configuration (inner level) along the axial direction is zero. The optimization is performed with Brent’s algorithm [74, 75] until changes in NT length are smaller than $10^{-12}$. After the optimization, all stress tensor components become zero.
According to the results of these numerical simulations, shown in Fig. 9 by continuous lines, the radius $R^{\prime}_{\mathrm{NT}}$ (length $L^{\prime}_{\mathrm{NT}}$) of the relaxed hBN nanotubes is slightly larger (smaller) than its approximate value given in Eq. (29) (Eq. (30)). The effect is slightly more pronounced in ZZ nanotubes. In any case for larger radius NTs, when $R_{\mathrm{NT}}>20$ , Eqs. (29) and (30) accurately represent the actual values.
Figure 9: Ratio of (a) the relaxed nanotube radius $R^{\prime}_{\textrm{NT}}$ over the approximate value of Eq. (29) and (b) the relaxed nanotube length $L^{\prime}_{\textrm{NT}}$ over the length of its planar counterpart given in Eq. (30). Red and blue colors correspond to ZZ and AC nanotubes, respectively. The circles and squares connected by solid lines represent evaluations using the original potential. Fine-dotted and dashed lines denote evaluations using a modified potential where the strength of the torsional potential terms (Eqs. (10) and (11)) is scaled by a factor of 0.1 and 10, respectively. Large dots illustrate evaluations with a modified potential where the strength of the bond-angle-bending term (Eq. (8)) is scaled by 0.1.
@@
The total deformation energy is dominated by the torsional terms which rise abruptly when the radius is decreasing, as expected. The contribution of bond-bending angles is non-negligible only at very small radii, affecting the structural relaxation illustrated in Fig. 9. Bonds do not contribute to the potential energy of relaxed NTs.
Figure 11: Variation of the relative surface energy per atom $\Delta\gamma_{\mathrm{NT}}$ (Eq. (32)), scaled by the factor ${R_{\mathrm{NT}}}^{2}$, with the NT radius for ZZ (circles) and AC (squares) nanotubes. Dashed line represents the analytically obtained bending rigidity ($D_{\textrm{sheet}}=1.563$ eV) from Eq. (34).
@@
(compare with Eqs. (6) and (30) in Ref. 77). Evaluating the first derivative of Eq. (4) at $\phi=2\pi/3$ and the second derivatives of Eqs. (10) and (11) at $\omega=0$ and $\omega=\pi$, respectively, Eq. (33) yields:
@@
$D_{\textrm{G,sheet}}=-\frac{8}{\sqrt{3}}(4k_{\mathrm{t}}^{\mathrm{cis}}+4k_{\mathrm{t}}^{\mathrm{trans}})\,.$ (36)
@@ III.3 Elastic Energy of hBN Nanotori and the Bending Rigidity of hBN Nanotubes
Figure 13: Total energy per atom of ZZ and AC hBN nanotori relative to that of the corresponding hBN nanotubes with the same internal radius and chirality, as a function of the torus external radius. Different colors represent nanotori of different internal radii, as indicated in the plots.
@@
Figure 14: Relative surface energy per atom $\Delta\gamma_{\mathrm{tor}}$, Eq. (40), multiplied by $2{R_{\mathrm{tor}}}^{2}$ for (a) ZZ and (b) AC hBN nanotori, as a function of the torus external radius. Different colors represent nanotori of different internal radii, as indicated in the plots. Dashed lines are guides to the eye and represent evaluations of the bending rigidity of the hBN nanotubes using Eq. (45).
Similarly to the discussion of the previous section, one can determine the bending rigidity of hBN nanotubes with different radii ($R_{\textrm{NT}}$) and chiralities, by considering the corresponding tori in the limit of large external radius through the relation:
$D_{\textrm{NT}}(R_{\textrm{NT}})\equiv\frac{\partial^{2}\gamma_{\mathrm{NT}}}{\partial\kappa^{2}}=\lim_{R_{\mathrm{tor}}\to\infty}2{R_{\mathrm{tor}}}^{2}\Delta\gamma_{\mathrm{tor}}$ (39)
@@
with in energy×$\times$×length units.
@@
According to Ref. [78], the elastic strain energy for bending a hollow cylinder (thickness $l_{0}$, radius $R_{\mathrm{NT}}$) into a ring of radius $R_{\mathrm{tor}}$ is
$E_{\textrm{tor}}-E_{\textrm{NT}}=\frac{{\pi}^{2}YR_{\mathrm{NT}}}{4R_{\mathrm{tor}}}\left((2R_{\mathrm{NT}})^{2}+{l_{0}}^{2}\right)\,,$ (42)
where we have adopted the notation in the present paper. By dividing Eq. (42) with the area of a torus, $S_{\textrm{tor}}~{}=~{}4\pi^{2}R_{\mathrm{NT}}R_{\mathrm{tor}}$, we can quantify the variation of the surface energy as follows:
@@
The thickness $l_{0}$ of the monolayer hBN sheet is in general much smaller than the torus’ external radius and thus the last term can be ignored, especially in the limit $R_{\mathrm{tor}}\to\infty$. Therefore, in this limit Eq. (43) reads
@@
Evaluations of Eq. (45) are illustrated with dashed lines in Fig. 14, for each value of the internal torus radii considered therein. In the limit of large $R_{\mathrm{rot}}$, the numerical evaluations of the bending rigidity ($2{R_{\mathrm{tor}}}^{2}\Delta\gamma_{\mathrm{tor}}$) are in excellent agreement with the analytical estimation from Eq. (45).
@@ Conclusions
We develop an atomistic force field for hexagonal boron nitride nanostructures including monolayers, nanotubes, and nanotori. The force field is parameterized via a systematic bottom-up procedure calibrated through first principle’s calculations (DFT) of properly deformed hBN monolayers, ensuring accurate representation of both in-plane and out-of-plane deformations.
The components of the stiffness matrix and the corresponding elastic constants (i.e., Young’s modulus, Poisson’s ratio, and bulk modulus) were calculated via a finite difference scheme using molecular mechanics simulations and validated against analytical formulas. The predicted mechanical properties are in agreement with experimental and theoretical values reported in the literature. For the potential energy terms considered in our force field model, the Poisson’s ratio of the nanosheets depends strictly on the relative strength of the bond-stretching and angle-bending interactions.
@@
By folding hBN nanotubes along their axial direction we generated hBN nanotori for a broad range of internal and external torus radii. The bending rigidity of hBN nanotubes with a particular radius was estimated through the deformation energy of the corresponding nanotori with the same internal radius in the limit of very large external radii. We found that above a certain threshold of NT radius, the nanotube bending rigidity increases quadratically with its radius.
The developed force field is simple, numerically efficient, yet effective, capable of accurately capturing the mechanical properties of hBN-based nanostructures. Its versatility could be useful for the development of hBN-based materials in applications such as thermal management, optoelectronics, and energy storage, among others.

### v2

@@ III.2 Elastic Energy of single layer hBN Nanotubes and the Bending Rigidity of hBN Monolayers
which is equivalent to the common expression $\frac{a}{2\pi}\sqrt{n^{2}+nm+m^{2}}$ [78] involving pairs $(n,m)$ of chiral indices, upon setting $(n,m)=(n_{\mathrm{ZZ}},0)$ for ZZ and $n=m=n_{\mathrm{AC}}$ for AC nanotubes. The initial length of the hBN nanotube is
@@
    The inner level (energy minimization) entails minimizing the potential energy of a NT with a fixed (input) length and returning the stress along the axial direction (output). The minimization is performed sequentially using the Polak-Ribiere conjugate-gradient [79] and the Hessian-free truncated Newton algorithms [65] until the norm of the global force vector is less than $10^{-9}\textrm{ ev}/$Å.
@@
    In the outer level (NT length optimization) the length of the NT is optimized until the stress of the minimized configuration (inner level) along the axial direction is zero. The optimization is performed with Brent’s algorithm [80, 81] until changes in NT length are smaller than $10^{-12}$. After the optimization, all stress tensor components become zero.
According to the results of these numerical simulations, shown in Fig. 9 by symbols connected with continuous lines, the radius $R^{\prime}_{\mathrm{NT}}$ (length $L^{\prime}_{\mathrm{NT}}$) of the relaxed hBN nanotubes is slightly larger (smaller) than its approximate value given in Eq. (29) (Eq. (30)). The effect is slightly more pronounced in ZZ nanotubes. In any case for larger radius NTs, when $R_{\mathrm{NT}}>20$ , Eqs. (29) and (30) accurately represent the actual values.
Figure 9: Ratio of (a) the relaxed nanotube radius $R^{\prime}_{\textrm{NT}}$ over the approximate value of Eq. (29) and (b) the relaxed nanotube length $L^{\prime}_{\textrm{NT}}$ over the length of its planar counterpart given in Eq. (30), as a function of the initial NT radius. Red and blue colors correspond to ZZ and AC nanotubes, respectively. The circles and squares connected by solid lines represent evaluations using the original potential. Fine-dotted and dashed lines denote evaluations using a modified potential where the strength of the torsional potential terms (Eqs. (10) and (11)) is scaled by a factor of 0.1 and 10, respectively. Large dots (overlapping with the dashed lines) illustrate evaluations with a modified potential where the strength of the bond-angle-bending term (Eq. (6)) is scaled by 0.1.
@@
The total deformation energy is dominated by the torsional terms which rise abruptly when the radius is decreasing, as expected. The contribution of bond-angle bending term is non-negligible only at very small radii. Bonds do not contribute to the potential energy of relaxed NTs.
Figure 11: Variation of the relative surface energy density $\Delta\gamma_{\mathrm{NT}}$ (Eq. (32)), scaled by the factor ${R_{\mathrm{NT}}}^{2}$, with the NT radius for ZZ (circles) and AC (squares) nanotubes. Dashed line represents the analytically obtained bending rigidity ($D_{\textrm{sheet}}=1.563$ eV) from Eq. (34).
@@
(compare with Eqs. (6) and (30) in Ref. 52). Evaluating the first derivative of Eq. (6) at $\phi=2\pi/3$ and the second derivatives of Eqs. (10) and (11) at $\omega=0$ and $\omega=\pi$, respectively, Eq. (33) yields:
@@
$D_{\textrm{G,sheet}}=-\frac{8}{\sqrt{3}}(4k_{\mathrm{t}}^{\mathrm{cis}}+4k_{\mathrm{t}}^{\mathrm{trans}})\,,$ (36)
@@ III.3 Elastic Energy of hBN Nanotori and the Bending Rigidity of hBN Nanotubes
Figure 13: Total energy per atom of (a) ZZ and (b) AC hBN nanotori relative to that of the corresponding hBN nanotubes with the same internal radius and chirality, as a function of the torus external radius. Different colors represent nanotori of different internal radii, as indicated in the plots.
@@
Figure 14: Relative surface energy density $\Delta\gamma_{\mathrm{tor}}$, Eq. (40), multiplied by $2{R_{\mathrm{tor}}}^{2}$ for (a) ZZ and (b) AC hBN nanotori, as a function of the torus external radius. Different colors represent nanotori of different internal radii, as indicated in the plots. Dashed lines represent evaluations of the bending rigidity of the hBN nanotubes using Eq. (45).
Similarly to the discussion of the previous subsection, one can determine the bending rigidity of hBN nanotubes with different radii ($R_{\textrm{NT}}$) and chiralities, by considering the corresponding tori in the limit of large external radius through the relation:
$D_{\textrm{NT}}(R_{\textrm{NT}})\equiv\frac{\partial^{2}\gamma_{\mathrm{NT}}}{\partial\kappa^{2}}=\lim_{R_{\mathrm{tor}}\to\infty}2{R_{\mathrm{tor}}}^{2}\Delta\gamma_{\mathrm{tor}}\,,$ (39)
@@
with energy×$\times$×length units.
@@
According to Ref. [83], the elastic strain energy for bending a hollow cylinder (thickness $l_{\mathrm{t}}$, radius $R_{\mathrm{NT}}$) into a ring of radius $R_{\mathrm{tor}}$ is
$E_{\textrm{tor}}-E_{\textrm{NT}}=\frac{{\pi}^{2}YR_{\mathrm{NT}}}{4R_{\mathrm{tor}}}\left((2R_{\mathrm{NT}})^{2}+{l_{\mathrm{t}}}^{2}\right)\,,$ (42)
where we have adapted the notation of the present paper. By dividing Eq. (42) with the area of a torus, $S_{\textrm{tor}}~{}=~{}4\pi^{2}R_{\mathrm{NT}}R_{\mathrm{tor}}$, we can quantify the variation of the relative surface energy density as follows:
@@
The corresponding thickness $l_{\mathrm{t}}$ of the monolayer hBN sheet can be approximatelly ignored, therefore, in this limit Eq. (43) reads
@@
Evaluations of Eq. (45) are illustrated with dashed lines in Fig. 14, for each value of the internal torus radii considered therein. In the limit of large $R_{\mathrm{rot}}$, the numerical evaluations of the bending rigidity ($2{R_{\mathrm{tor}}}^{2}\Delta\gamma_{\mathrm{tor}}$) are in excellent agreement with the analytical estimation from Eq. (45). Surprisingly, the variation of $2{R_{\mathrm{tor}}}^{2}\Delta\gamma_{\mathrm{tor}}$ across the stable $R_{\mathrm{tor}}$ regime does not monotonically approach its limiting value corresponding to the NT bending rigidity. Instead, it exhibits a maximum as $R_{\mathrm{tor}}$ increases before asymptotically approaching the limiting constant from above.
@@ Conclusions
We develop an atomistic force field for hexagonal boron nitride nanostructures including 2D monolayers, single-layer nanotubes, and nanotori. The force field is parameterized via a systematic bottom-up procedure calibrated through first principle’s calculations (DFT) of properly deformed hBN monolayers, ensuring accurate representation of both in-plane and out-of-plane deformations. The parameterization included data from significantly deformed sheets accounting for nonlinear effects, but it did not extend to extreme deformations. Additional density functional theory calculations, performed using Gaussian16[85] on hydrogen-passivated BN flakes (results not shown), revealed that the electronic character of hBN remains largely unaffected by out-of-plane bending. This stability is attributed to the robust electronic structure, as evidenced by the large HOMO-LUMO gap.
The components of the stiffness matrix and the corresponding elastic constants (i.e., Young’s modulus, Poisson’s ratio, and bulk modulus) were calculated via a finite difference scheme using molecular mechanics calculations and validated against analytical formulas. The predicted mechanical properties are in agreement with experimental and theoretical values reported in the literature. For the potential energy terms considered in our force field model, the Poisson’s ratio of the nanosheets depends strictly on the relative strength of the bond-stretching and angle-bending interactions.
@@
The elastic energy of carbon nanotori has been successfully described using a continuum model for bent hollow beams [83]. Building on this framework, by applying the fundamental definition of bending rigidity to the continuum model we derived a model for predicting the bending rigidity of nanotubes. To assess the accuracy of our model, we compared its predictions with numerical estimations of bending rigidity of hBN nanotubes with a given radius, obtained through the deformation energy of the corresponding nanotori with the same internal radius in the limit of very large external radii. We found that the nanotube bending rigidity increases quadratically with its radius and proportionally with Young’s modulus, in agreement with our model. Interestingly, while the relative surface energy density follows a ${R_{\mathrm{tor}}}^{-2}$ scaling at large radii, its behavior becomes non-monotonic for smaller radii above the critical buckling point. Within this regime, $2{R_{\mathrm{tor}}}^{2}\Delta\gamma_{\mathrm{tor}}$ reaches a maximum with increasing ${R_{\mathrm{tor}}}$ before asymptotically approaching its limiting value (associated with nanotube bending rigidity) from above.
The proposed force field belongs to the class of classical, non-reactive force fields, sharing the advantages of simplicity, ease of use, and efficiency, similar to relevant models in the literature [73, 71, 70, 74]. A key distinction of our implementation, however, lies in the incorporation of nonlinear terms in the in-plane bond-stretching and bond-angle-bending potentials. These terms account for anharmonic effects, which become significant at large deformations and/or elevated temperatures.
As demonstrated in Ref [41] through a direct comparison between nonlinear and linearized potentials for graphene, the inclusion of the nonlinear terms considerably improves the accuracy of the predicted elastic response under large uniaxial extensions; see Fig. 7, therein. The linearized potentials become very inaccurate beyond strain values of 0.05 where the anharmonic effects become important, Poisson’s ratio is particularly sensitive to anharmonic effects as its qualitative behavior changes significantly in their absence [41]. In addition, linearized force fields fail to capture phonon softening—the redshift of the phonon density of states—that occurs with increasing strain due to weakened interatomic interactions.
Tersoff-based potentials [32, 76, 77, 30, 31] and their extensions (e.g., the ExTeP force field [34]), also employ nonlinear functions to model interatomic interactions, similar to our method. This enables them to qualitatively capture the elastic behavior of h-BN under large mechanical strains as well. However, the reactive and bond-order character of these potentials requires continuous monitoring of the local coordination number and potential reactions. These processes involve constructing neighboring lists during the simulations, which increases the communication overhead and makes them computationally expensive. Indeed, according to benchmarks (results not shown in the Arxiv version), the proposed force field is roughly five times faster than a Tersoff-based potential [34].
An additional utility of the class of nonreactive force fields lies in the ease with which their parameters can be related to the resulting model behavior. Specifically, simple closed-form expressions enable the exact calculation of both in-plane and out-of-plane elastic constants at 0 K through Eqs. (24)-(26) and (33)-(35), respectively. These analytical expressions are generalizable to any 2D membrane model that shares the same crystalline structure and consists solely of bonds, bond angles, and dihedral angles. This feature is particularly valuable for top-down parameterization of such 2D membranes, especially in mesoscopic coarse-grained modeling. For example, if the in-plane and out-of-plane elastic constants are known, the corresponding force field parameters can be determined analytically, ensuring the reproduction of the target elastic properties at 0 K.
To summarize, the developed force field is simple, numerically efficient, yet effective, capable of accurately capturing the mechanical properties of hBN nanostructures in both equilibrium and nonlinear regimes. Its versatility could be useful for the development of hBN-based materials in applications such as thermal management, optoelectronics, and energy storage, among others.
Data Availability Statement
The data that support the findings of this article are openly available [86].